OpenStudy (dominirican1013):

Solve on the interval [0,2π): 2cos2x+3cosx+1=0

OpenStudy (michele_laino):

here we have to make this substitution: \[\cos \left( {2x} \right) = 2{\left( {\cos x} \right)^2} - 1\]

OpenStudy (michele_laino):

so your equation, can be rewritten as follows: \[2\left( {2{{\left( {\cos x} \right)}^2} - 1} \right) + 3\cos x + 1 = 0\]

OpenStudy (michele_laino):

please simplify

OpenStudy (michele_laino):

what equation do you get?

OpenStudy (dominirican1013):

I don't know I'm trying to work it out

OpenStudy (michele_laino):

hint: \[4{\left( {\cos x} \right)^2} - 2 + 3\cos x + 1 = 0\]

OpenStudy (dominirican1013):

I'll be right back.

OpenStudy (dominirican1013):

I don't know how to do this.

OpenStudy (michele_laino):

more explanation: we have: \[2\left( {2{{\left( {\cos x} \right)}^2} - 1} \right) = 4{\left( {\cos x} \right)^2} - 2\] am I right?

OpenStudy (dominirican1013):

these are the choices \[A. x=2\pi, x=\pi/3\] \[B. x=\pi, x=2\pi/3, x=4\pi/3\] \[C. x=2\pi, x=\pi/4, x=5\pi/4\] \[D. x=\pi/6, x=7\pi/6\]

OpenStudy (michele_laino):

yes! I see them, nevertheless I can not give you the answer directly, since I have to respect the Code of Conduct

OpenStudy (michele_laino):

If I simplify my equation above, I get: \[4{\left( {\cos x} \right)^2} + 3\cos x - 1 = 0\] now, I make this substitution: cos(x)=y, so I can write: \[4{y^2} + 3y - 1 = 0\] which is a quadratic equation. please solve that equation for y

OpenStudy (dominirican1013):

B is the answer I got it